solve the equation:
6[x-(2x+3)]=5(1-x)-1
thanks a lot

Answer by checkley71(8403)

6[X-(2X+3)]=5(1-X)-1 OR 6[X-2X-3]=5-5X OR 6[-X-3]=5-5X-1 OR -6X-18=4-5X OR
-X=22 OR X=-22